# Attention-Based Beamformer For Multi-Channel Speech Enhancement

### v1

@@ IV-B Ablation experiments
    ONLINE-MVDR / BLOCK-MVDR: To validate the facilitation of the ISAM module, this ablation experiment calculates the SCM using Adaptive-based [7] (ONLINE-MVDR) and Block-based [8] (BLOCK-MVDR) based on the ISCM. The forgetting factor is set to 0.995 in ONLINE-MVDR and the block size is set to 30 in BLOCK-MVDR.
@@
Table III shows the results of the ablation experiments on moving datasets. Compared to ONLINE-MVDR and BLOCK-MVDR, the proposed model demonstrates superior performance. This suggests that the inclusion of the ISAM module enhances the capability of the SCM reconstruction and target speaker tracking in complex scenarios. From Table III and IV, it can be seen that the proposed model outperforms CRN-MVDR, with reduced computation and fewer parameters. This proves that spatial information is effectively extracted through IGCRN for reconstructing SCMs, making it compatible with beamformer.
@@ V Conclusion
This paper introduces a novel approach that combines IGCRN and an attention-based mechanism to estimate speech and noise SCM, followed by applying MVDR to obtain enhanced speech. The IGCRN leverages the characteristics of the inplace convolution operator and frequency-independent LSTM to learn spatial information. The attention-based mechanism is applied to allocate temporal weights for reconstructing the speech and noise SCMs accurately. The experimental results demonstrate that the proposed method performs baselines under various conditions. Furthermore, due to the small number of parameters, the proposed model can be easily deployed in edge devices.

### v2

@@ IV-B Ablation experiments
    ONLINE-MVDR / BLOCK-MVDR: To validate the facilitation of the ISAM module, this ablation experiment calculates the SCM using online processing [7, 9] (ONLINE-MVDR) and blockwise processing [8] (BLOCK-MVDR) based on the ISCM. The forgetting factor is set to 0.995 in ONLINE-MVDR and the block size is set to 30 in BLOCK-MVDR.
@@
Table III shows the results of the ablation experiments on moving datasets. Compared to ONLINE-MVDR and BLOCK-MVDR, the proposed model demonstrates superior performance. This suggests that the inclusion of the ISAM module enhances the capability of SCM reconstruction, allowing it to allocate weights more accurately over time compared to online or blockwise processing in scenarios with moving sources. From Table III and IV, it can be seen that the proposed method outperforms CRN-MVDR on most metrics, with reduced computation and fewer parameters. This proves that spatial information is effectively extracted through IGCRN for reconstructing SCMs, making it compatible with beamformer.
@@ V Conclusion
This paper proposes a novel attention-based beamformer combining IGCRN for multi-channel speech enhancement. The IGCRN leverages the characteristics of the inplace convolution operator and frequency-independent LSTM to learn spatial information. The attention mechanism is applied to allocate temporal weights for reconstructing the speech and noise SCMs accurately. The MVDR beamformer reduces speech distortion. Experimental results indicate that the proposed method surpasses the baselines across various conditions, with reduced computational complexity and fewer parameters.